# Computing Lower and Upper Hitting Probabilities for Imprecise Markov Chains

### v1

@@ 6.1 Description of the Experimental Setup
First, we explain the experimental setup. Given $N$ the number of states of $\mathcal{X}$, we set $A=\{1\}$ to be the target and construct a random graph with $N$ nodes where each directed edge indicates a positive upper probability of transition between the corresponding states. We construct the random graph in the following steps.
@@
    State $N$ is made absorbing, i.e. any random walk that reaches $N$ remains there with probability one. We impose this condition so that the upper hitting probability may be less than one.
@@
    We construct a random tree rooted in the target set $A$. Specifically, let $\pi$ be a random permutation of $\mathcal{X}\setminus(A\cup\{N\})$; for each state $x$ in the order given by $\pi$, we draw a directed edge from $x$ to either $A$ or to a state chosen uniformly from the set of states that have already been connected to the target set. This construction is chosen so that every state but $N$ has upper hitting probability greater than zero.
@@
    We connect any two nodes of $\mathcal{X}\setminus(A\cup\{N\})$ with probability $\nicefrac{{(\lambda-1)}}{{(N-1)}}$, where $1\leq\lambda\leq N$. This Erdős-Rényi-like [erdos1960evolution] construction ensures that each node has an average degree of $\lambda$. In this way, we can easily control the graph’s density and study its influence on the average number of iterations.
We now generate the set of transition matrices $\mathcal{T}$. To do that, we propose the following two methods:
@@
    Given a parameter $0<\varepsilon<1$, we sample a transition matrix $\tilde{T}$ uniformly at random and $\varepsilon$-contaminate it [Huber1964, itip:specialcases]. In particular, the set of transition matrices is T={(1-ε)~T+εS ∣ for all trans. mat SS}. For numerical experiments, we set $\varepsilon=0.1$; experiments with other values exhibited a similar behaviour.
@@
    We sample $n_{x}$ points uniformly at random from the $n_{x}-1$-dimensional simplex, where $n_{x}$ is the number of outgoing edges of $x\in\mathcal{X}$, and $\mathcal{T}_{x}$ is their convex hull. Since the set of transition matrices $\mathcal{T}$ has separately specified rows, $\mathcal{T}$ is the Cartesian product of all $\mathcal{T}_{x}$.
@@ 6.2 Numerical Results
Since the algorithms from Propositions 4.1 and 4.3 exhibit nearly identical iteration counts, we will often present the number of iterations without explicitly specifying which algorithm it refers to. Figure 2 reports the average number of iterations, on the vertical axis, as a function of the average degree $\lambda$ of the graph, on the horizontal axis. The graphs contain either 10 (left) or 50 (right) nodes, and the set of transition matrices is generated using $\varepsilon$-contamination (top) or by taking the convex hull of random samples (bottom). The averages are obtained by running 1000 independent simulations.
Figure 2: The red and blue lines represent the average number of iterations needed to compute lower and upper hitting probabilities, respectively, as a function of the average degree of each node of the graph, $\lambda$. The two magenta shaded regions correspond to the intervals within one standard deviation ($\pm\sigma$) and approximately $95\%$ confidence ($\pm 1.96\sigma$) around the mean number of iterations.
@@
Figure 3: Number of instances, out of 1000, on which the algorithm for upper hitting probabilities converged in a given number of iterations, for different values of the average degree $\lambda$. The graph has $N=50$ nodes and the set of transition matrices is generated by taking the convex hull of transition matrices sampled uniformly at random.
Furthermore, we note that the number of iterations for the complete graph with $N$ nodes and set of transition matrices generated via $\varepsilon$-contamination is always two. This is a consequence of the fact that the credal sets of this uncertainty model are probability intervals of width $\varepsilon$, namely
@@
for all $y\in A^{c}$. The second and last iteration computes the upper hitting probability by solving the linear system and checks that no further improvement can be made. The same happens with lower hitting probabilities with $\{N\}$ in place of $\{1\}$.
@@
Figure 4: Average number of iterations as a function of the number of states $N$ and the average degree $\lambda$. Each red dot on the blue line marks, for a fixed number of states $N$, the average degree $\lambda^{*}$ that attains the maximum average number of iterations.
@@
$\mathcal{T}=\left\{T=\begin{bmatrix}1&0&0&\dots&\dots&\dots&0\\ q_{2}&0&0&\dots&\dots&\dots&1-q_{2}\\ 0&q_{3}&0&\dots&\dots&1-q_{3}&0\\ 0&0&q_{4}&0&\dots&1-q_{4}&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\dots&\dots&0&q_{N}&1-q_{N}&0\\ \nicefrac{{1}}{{2}}&0&\dots&\dots&\dots&0&\nicefrac{{1}}{{2}}\\ 0&\dots&\dots&\dots&\dots&0&1\end{bmatrix}:\begin{aligned} &\;\;q_{x}\in[0,b],\\ &\sqrt[N]{\nicefrac{{1}}{{2}}}<b<1\end{aligned}\right\}.$
@@
Initialise the algorithm to compute upper hitting probabilities with any matrix $T_{1}$ satisfying $0<q_{2}<\nicefrac{{1}}{{2}}$. Then, $p_{1}(2)<\nicefrac{{1}}{{2}}$ while trivially $p_{1}(1)=1$, $p_{1}(N+2)=0$, $p_{1}(N+1)=\nicefrac{{1}}{{2}}$, and $p_{1}(x)\leq\nicefrac{{1}}{{2}}$ for every other $x\in\mathcal{X}$.
In the second iteration, node $3$ sets $q_{3}=0$ because it is connected to nodes $2$ and $N+1$, and $b\cdot p_{1}(2)<p_{1}(N+1)=\nicefrac{{1}}{{2}}$. The same argument applies for every other node $x\in\{4,\dots,N\}$: $q_{x}$ is set to zero because $b\cdot p_{1}(x-1)<p_{1}(N+1)=\nicefrac{{1}}{{2}}$. Meanwhile, node $2$ assigns probability $b$ to the transition to the target set $\{1\}$. Henceforth, solving the linear system yields $p_{2}(2)=b$ and $p_{2}(x)=\nicefrac{{1}}{{2}}$, for all $x\in\{3,\dots,N\}$.
In the third iteration, we have $b^{2}>p_{2}(N+1)=\nicefrac{{1}}{{2}}$, so node $3$ now sets $q_{3}=b$, yielding a hitting probability of $p_{3}(3)=b^{2}$. The information that larger $q_{x}$ yields greater hitting probabilities propagates backward one node per iteration, so that only after $N$ iterations the algorithm has reached convergence.
@@ 7 Conclusions and Future Work
In Section 3.1 we have discussed in detail the notion of reachability in the imprecise setting. The literature review highlighted several different definitions of lower reachability; hence we have explored the connections among them, presenting examples to clarify their logical relations. In particular, we have shown the equivalence between definitions $\operatorname{(LR0)}$ and $\operatorname{(LR1)}$; this equivalence has proved crucial, as the latter definition provides a practical algorithm to identify the states that cannot lower reach the target and thus have zero lower hitting probability. Using the weakest form of lower reachability, $\operatorname{(LR0)}$, we have been able to partition the state space and characterise lower hitting probabilities as the unique solution of a nonlinear fixed-point equation. An analogous characterisation has been proved for upper hitting probabilities, as they too are the unique solution to a similar fixed-point equation.
In Section 4 we have presented novel efficient iterative algorithms that compute lower and upper hitting probabilities. These algorithms alternate between solving a linear system over nontrivial states and selecting extreme points from the set of transition matrices. Although in principle these algorithms may visit all extreme points, this has never occurred in practice during numerical experiments. In fact, our experiments, described in Section 6, have shown the efficiency of the algorithms: for every graph structure and uncertainty model analysed, very few iterations were required for convergence.
@@
    Formal justification of the nonmonotone relation between density and number of iterations. We would like to investigate why the algorithms for computing lower and upper hitting probabilities show a nonmonotone relation between the density $\lambda$ of the graph and the average number of iterations. For specific uncertainty models (e.g. $\varepsilon$-contamination, probability intervals, etc.), we may try to find the average degree $\lambda^{*}$ that maximises the number of iterations.
@@
    Handling pathological credal sets. Although extremely rare, some credal set shapes force the algorithms to cycle over all extreme points, as in Example 5.3. Future work should identify such cases in advance and aim to derive specialised algorithms to compute lower and upper hitting probabilities, that perform better in such regimes.
@@
    Extension to continuous-time IMCs. One relatively straightforward step forward, would be to generalise the fixed-point characterisation and the computational framework discussed here, to continuous-time imprecise Markov chains.
@@
    Lower and upper meeting probabilities. It would be interesting to characterise and compute efficiently lower and upper meeting probabilities, i.e. the probability that two (or more) processes are ever in the same state at the same time. This present paper, combined with Sangalli et al.’s recent work on meeting times for imprecise Markov chains [sangalli2025meeting], provides a foundation for this line of research.
@@
    Relation between lower and upper hitting times and probabilities. Finally, we may aim to investigate the relation between hitting times and probabilities in the imprecise framework, and seek to establish whether knowledge of one quantity allows us to infer bounds on the other.
@@ Acknowledgements
This work has been partly supported by the PersOn project (P21-03), which has received funding from Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO).
@@ Appendix A Proofs of Auxiliary Lemmas
We now check that $y\in A^{c}\setminus\mathcal{C}_{T}$. First, $y\notin A$: if $y\in A$ then $T(x,y)>0$ would imply $T\mathbb{1}_{A}](x)\geq T(x,y)>0$, contradicting the minimality assumption that $T^{m}\mathbb{1}_{A}](x)=0$ for every $m<n$ (recall $n>1$). Thus $y\in A^{c}$.
Second, $y\notin\mathcal{C}_{T}$: by definition any $z\in\mathcal{C}_{T}$ satisfies $T^{m}\mathbb{1}_{A}](z)=0$ for all $m\in\mathbb{N}$, but we have $T^{\,n-1}\mathbb{1}_{A}](y)>0$, so $y\notin\mathcal{C}_{T}$.
Therefore $y\in A^{c}\setminus\mathcal{C}_{T}$, $T(x,y)>0$ and $T^{\,n-1}\mathbb{1}_{A}](y)>0$, which proves the lemma. ∎
@@
But $M^{k}\mathbf{0}=\mathbf{0}$ for every $k$. Hence $\mathbf{0}\leq M^{k}\mathbf{1}$.
For the upper bound, observe that for every $x\in\mathcal{S}$
@@
Combining the two inequalities gives $\mathbf{0}\leq M^{k}\mathbf{1}\leq\mathbf{1}$ for all $k\in\mathbb{N}_{0}$, which completes the proof. ∎
@@
Consider any path of length $n_{x}$, $x_{0},\dots,x_{n_{x}}$, with $x_{0}=x$ and $x_{n_{x}}\in A$. Because $m<n_{x}$ implies $T^{m}\mathbb{1}_{A}](x)=0$, such a path cannot visit $A$ before time $n_{x}$. Moreover, by the definition of $\mathcal{C}_{T}$, a path that ever enters $\mathcal{C}_{T}$ cannot reach $A$ at any later time; hence every path contributing to $T^{n_{x}}\mathbb{1}_{A}](x)$ stays in $A^{c}\setminus\mathcal{C}_{T}$} up to time $n_{x}-1$ and then jumps from $A^{c}\setminus\mathcal{C}_{T}$} to $A$ at time $n_{x}$.
@@
Now we use Lemma 2.8: since $M{\mathbf{1}}\leq{\mathbf{1}}$, the map $k\mapsto M^{k}{\mathbf{1}}$ is componentwise nonincreasing: $M^{k+1}{\mathbf{1}}=M^{k}(M{\mathbf{1}})\leq M^{k}{\mathbf{1}}$ for all $k\in\mathbb{N}_{0}$. Let $n\coloneqq\max_{x\in{A^{c}\setminus\mathcal{C}_{T}}}n_{x}$, which is finite because $A^{c}\setminus\mathcal{C}_{T}$} is finite. Then for any $k\geq n$ and any $x\in{A^{c}\setminus\mathcal{C}_{T}}$,
$M^{k}{\mathbf{1}}](x)\leq[M^{n_{x}}{\mathbf{1}}](x)<1.$
This proves the claim. ∎
@@
We show by induction that $(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}](x)\leq[(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}](x)$ for all $x\in A^{c}\setminus\mathcal{S}$ and all $k\in\mathbb{N}$.
First consider $k=1$. Then
$[T|_{A^{c}\setminus\mathcal{S}}\mathbf{1}](x)$ $=[T\mathbb{1}_{A^{c}\setminus\mathcal{S}}](x)$
$\leq[T\mathbb{1}_{A^{c}\setminus\mathcal{C}_{T}}](x)$
$=[T|_{A^{c}\setminus\mathcal{C}_{T}}\mathbf{1}](x),$
where the inequality follows from the linearity of $T$ and the fact that we assumed $\mathcal{C}_{T}\subseteq\mathcal{S}$.
Now let us assume as induction hypothesis that
$\Bigl[(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}\Bigr](y)\leq\Bigl[(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}\Bigr](y)$
for all $y\in A^{c}\setminus\mathcal{S}$ for some $k\in\mathbb{N}$. We will show that the same statement then also holds for $k+1$. We find that
$\Bigl[(T|_{A^{c}\setminus\mathcal{S}})^{k+1}\mathbf{1}\Bigr](x)$ $=[T|_{A^{c}\setminus\mathcal{S}}(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}](x)$
$=\Bigl[T\Bigl(((T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](x)$
$=\sum_{y\in\mathcal{X}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y)$
$=\sum_{y\in A^{c}\setminus\mathcal{S}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y)$
$\leq\sum_{y\in A^{c}\setminus\mathcal{S}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y),$
where the inequality follows from the induction hypothesis. We now add the terms of $(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}$ corresponding to the states in $A^{c}\cap(\mathcal{S}\setminus\mathcal{C}_{T})$ to the sum, which are nonnegative due to Lemma 2.8. This then yields that
$\Bigl[(T|_{A^{c}\setminus\mathcal{S}})^{k+1}\mathbf{1}\Bigr](x)$ $\leq\sum_{y\in A^{c}\setminus\mathcal{C}_{T}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y)$
$=\sum_{y\in\mathcal{X}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y)$
$=\Bigl[T\Bigl(((T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](x)$
$=[T|_{A^{c}\setminus\mathcal{C}_{T}}\,(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}](x)$
$=\Bigl[(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k+1}\mathbf{1}\Bigr](x).$
This completes the induction proof. Hence we have shown that
$\Bigl[(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}\Bigr](x)\leq\Bigl[(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}\Bigr](x)$
for all $x\in A^{c}\setminus\mathcal{S}$ and all $k\in\mathbb{N}$. Applying Lemma 2.9 now yields the result. ∎
@@
Let $\mathcal{S}\subseteq\mathcal{X}$ satisfy $\mathcal{C}_{T}\subset\mathcal{S}\subset A^{c}$ and set
@@
$M^{n}\mathbf{1}=M^{n}(-\mathbf{1})\leq M^{n}f\leq M^{n}\mathbf{1}.$
[…]
which proves the claim. ∎

### v2

@@ 6.1 Description of the Experimental Setup
We explain the experimental setup. Given $N$ the number of states of $\mathcal{X}$, we set $A=\{1\}$ to be the target state. First, we construct a random graph with $N$ nodes, where each directed edge indicates a positive upper probability of transition between the corresponding states. Then, we use this graph to generate the set of transition matrices $\mathcal{T}$. We construct the random graph in the following three steps.
@@
    State $N$ is made absorbing, i.e. any random walk that reaches state $N$ remains there with probability one. We impose this condition so that the upper hitting probability can be less than one.
@@
    We construct a random tree rooted in $A=\{1\}$. Specifically, let $\pi$ be a random permutation of $\mathcal{X}\setminus(A\cup\{N\})$; for each state $x$ in the order given by $\pi$, we draw a directed edge from $x$ to a state chosen uniformly at random from the set $A\cup\{y\mid\pi(y)\leq\pi(x)\}$. This construction is chosen so that every state but $N$ has an upper hitting probability greater than zero.
@@
    For each pair of nodes, $x\in\mathcal{X}\setminus(A\cup\{N\})$ and $y\in\mathcal{X}$, we add a directed edge from $x$ to $y$ with probability $\nicefrac{{(\lambda-1)}}{{(N-1)}}$, where $1\leq\lambda\leq N$. This Erdős-Rényi-like [10] construction ensures that each node has an average degree of $\lambda$. In this way, we can easily control the graph’s density and study its influence on the average number of iterations.
We now generate the set of transition matrices $\mathcal{T}$. To do that, we consider the following two methods:
@@
    Given a parameter $0<\varepsilon<1$, we sample a transition matrix $\tilde{T}$ uniformly at random [^4] In particular, for each $x\in\mathcal{X}\setminus(A\cup\{N\})$, we sample a point $p_{x}$ uniformly at random from the $n_{x}-1$–dimensional simplex, and then we construct $\tilde{T}$ row-wise, i.e. $\tilde{T}|_{E_{x}}(x,\cdot)=p_{x}$, where $E_{x}$ is the set of states $x$ is connected to. and we $\varepsilon$-contaminate it [16, 7]. In particular, the set of transition matrices is
    $\mathcal{T}=\{(1-\varepsilon)\tilde{T}+\varepsilon S\mid\text{ for all trans. mat $S$}\}.$
    For numerical experiments, we set $\varepsilon=0.1$; experiments with other values exhibited a similar behaviour.
@@
    For each $x\in\mathcal{X}\setminus(A\cup\{N\})$, we sample $n_{x}$ points [^5] We choose this number of points to be sampled so that $\mathcal{T}_{x}$ has the same dimension as the simplex in which it is embedded. uniformly at random from the $n_{x}-1$-dimensional simplex, where $n_{x}$ is the number of outgoing edges of $x\in\mathcal{X}$, and $\mathcal{T}_{x}$ is their convex hull. Since the set of transition matrices $\mathcal{T}$ has separately specified rows, $\mathcal{T}$ is the Cartesian product of all $\mathcal{T}_{x}$.
@@ 6.2 Numerical Results
Since the algorithms from Propositions 4.1 and 4.3 exhibit nearly identical iteration counts, we will often present the number of iterations without explicitly specifying which algorithm it refers to. Figure 2 displays the average number of iterations, on the vertical axis, as a function of the average degree $\lambda$ of the graph, on the horizontal axis. The graphs contain either 10 (left) or 50 (right) nodes, and the set of transition matrices is generated using $\varepsilon$-contamination (top) or by taking the convex hull of random samples (bottom). The averages are obtained by running 1000 independent simulations.
Figure 2: The red and blue lines represent the average number of iterations needed to compute lower and upper hitting probabilities, respectively, as a function of the average degree $\lambda$ of each node of the graph. The two magenta shaded regions correspond to the intervals within one standard deviation ($\pm\sigma$) and approximately $95\%$ confidence ($\pm 1.96\sigma$) around the mean number of iterations.
@@
Figure 3: Number of instances, out of 1000, in which the algorithm for upper hitting probabilities converged in a given number of iterations, for different values of the average degree $\lambda$. The graph has $N=50$ nodes and the set of transition matrices is generated by taking the convex hull of transition matrices sampled uniformly at random.
Furthermore, the number of iterations for the complete graph with $N$ nodes and set of transition matrices generated via $\varepsilon$-contamination is always two. This is a consequence of the fact that the credal sets of this uncertainty model are probability intervals of width $\varepsilon$, namely
@@
for all $y\in A^{c}$. The second (and last) iteration computes the upper hitting probability by solving the linear system and checks that no further improvement can be made. The same happens with lower hitting probabilities with $\{N\}$ in place of $\{1\}$.
@@
Figure 4: Average number of iterations as a function of the number of states $N$ and the average degree $\lambda$. Each red dot on the blue line marks, for a fixed number of states NN, the average degree $\lambda^{*}=\lambda^{*}(N)$ that attains the maximum average number of iterations.
@@
$\mathcal{T}=\left\{T=\begin{bmatrix}1&0&0&\dots&\dots&\dots&0\\ q_{2}&0&0&\dots&\dots&\dots&1-q_{2}\\ 0&q_{3}&0&\dots&\dots&1-q_{3}&0\\ 0&0&q_{4}&0&\dots&1-q_{4}&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\dots&\dots&0&q_{N}&1-q_{N}&0\\ \nicefrac{{1}}{{2}}&0&\dots&\dots&\dots&0&\nicefrac{{1}}{{2}}\\ 0&\dots&\dots&\dots&\dots&0&1\end{bmatrix}:\begin{aligned} &\ q_{x}\in[0,b],\\ &{\nicefrac{{1}}{{2}}}<b<1\end{aligned}\right\}.$
@@
Initialise the algorithm to compute upper hitting probabilities with any matrix $T_{1}$ satisfying $0<q_{2}<\nicefrac{{1}}{{2}}$ and $q_{3}=b$. Then, $p_{1}(1)=1$, $p_{1}(N+2)=0$, $p_{1}(N+1)=\nicefrac{{1}}{{2}}$, $p_{1}(2)<\nicefrac{{1}}{{2}}$, $p_{1}(3)=bp_{1}(2)+(1-b)p_{1}(N+1)<\nicefrac{{1}}{{2}}$ and $p_{1}(x)\leq\nicefrac{{1}}{{2}}$ for every other $x\in\mathcal{X}$.
In the first iteration, node $2$ maximises its probability of hitting the target by setting $q_{2}=b$. Node $3$ sets $q_{3}=0$ because it is connected to nodes $2$ and $N+1$, and $b\cdot p_{1}(2)+(1-b)p_{1}(N+1)<p_{1}(N+1)$ since $p_{1}(2)<p_{1}(N+1)=\nicefrac{{1}}{{2}}$. The same argument applies for every other node $x\in\{4,\dots,N\}$: $q_{x}$ is set to zero because $b\cdot p_{1}(x-1)+(1-b)p_{1}(N+1)<p_{1}(N+1)$ which follows from $p_{1}(x-1)<p_{1}(N+1)=\nicefrac{{1}}{{2}}$. Henceforth, solving the linear system yields $p_{2}(2)=b$ and $p_{2}(x)=\nicefrac{{1}}{{2}}$, for all $x\in\{3,\dots,N\}$.
In the second iteration, we have $b\cdot p_{2}(2)+(1-b)p_{2}(N+1)>p_{2}(N+1)$ because $b=p_{2}(2)>p_{2}(N+1)=\nicefrac{{1}}{{2}}$, so node $3$ now sets $q_{3}=b$, yielding a hitting probability of $p_{3}(3)=b^{2}+\nicefrac{{1}}{{2}}(1-b)$. Other values of $q_{x}$ remain unchanged in this iteration.
In the $k$-th iteration, we have $b\cdot p_{k}(k)+(1-b)p_{k}(N+1)>p_{k}(N+1)=\nicefrac{{1}}{{2}}$ and $q_{k+1}$ is set to $b$. It can be shown (e.g. using induction) that $p_{k}(k)=\nicefrac{{1}}{{2}}+b^{k-2}(b-\nicefrac{{1}}{{2}})$ which is greater than $\nicefrac{{1}}{{2}}$ if $b>\nicefrac{{1}}{{2}}$.
Therefore, the information that larger $q_{x}$ yields greater hitting probabilities propagates backward one node per iteration, so that only after $N$ iterations the algorithm has reached convergence.
@@ 7 Conclusions and Future Work
In Section 3.1 we have discussed in detail the notion of reachability in the imprecise setting. The literature review highlighted several different definitions of lower reachability; hence we have explored the connections among them, presenting examples to clarify their logical relations. In particular, we have shown the equivalence between definitions $\operatorname{(LR0)}$ and $\operatorname{(LR1)}$; this equivalence has proved crucial, as the latter definition provides a practical algorithm to identify the states that cannot lower reach the target and thus have zero lower hitting probability. Using $\operatorname{(LR0)}$, we have been able to partition the state space and characterise lower hitting probabilities as the unique solution of a nonlinear fixed-point equation. The analogous characterisation for upper hitting probabilities fails; we needed to add stronger assumptions on the set $\mathcal{T}$ to restore uniqueness of the solution $\overline{p}^{\mathcal{T}}$.
In Section 4 we have presented novel efficient iterative algorithms that compute lower and upper hitting probabilities. These algorithms alternate between solving a linear system over nontrivial states and selecting extreme points from the set of transition matrices. Although in principle these algorithms may visit all extreme points of $\mathcal{T}$, this has never occurred in practice during numerical experiments. In fact, our experiments, described in Section 6, have shown the efficiency of the algorithms: for every graph structure and uncertainty model analysed, very few iterations were required for convergence.
@@
    Formal justification of the nonmonotone relation between density and number of iterations. We would like to investigate why the algorithms for computing lower and upper hitting probabilities show a nonmonotone relation between the density $\lambda$ of the graph and the average number of iterations. For specific uncertainty models (e.g. $\varepsilon$-contamination, probability intervals, etc.), we wish to find the average degree $\lambda^{*}=\lambda^{*}(N)$ that maximises the number of iterations.
@@
    Handling pathological credal sets. Although extremely rare, some credal set shapes force the algorithms to cycle over all extreme points, as in Example 5.3. Future work should identify such cases in advance and aim to derive specialised algorithms to compute lower and upper hitting probabilities that perform better in such regimes.
@@
    Extension to continuous-time IMCs. One relatively straightforward step forward would be to generalise the fixed-point characterisation and the computational framework discussed here to continuous-time imprecise Markov chains.
@@
    Algorithms for conservative inference. It follows from Corollary 5.1 that the proposed algorithms provide, at each iteration $n\in\mathbb{N}$ before convergence, an interval which is an inner approximation of the actual lower/upper hitting probabilities $\underline{p}^{\mathcal{T}},\overline{p}^{\mathcal{T}}]$, and which monotonically approaches these values. An interesting avenue of research consists of finding algorithms that instead compute a conservative approximation of lower and upper hitting probabilities, that is, an outer approximation that shrinks to these target values.
@@
    Lower and upper meeting probabilities. It would be interesting to characterise and compute efficiently lower and upper meeting probabilities, i.e. the probability that two (or more) processes are ever in the same state at the same time. This present paper, combined with Sangalli et al.’s recent work on meeting times for imprecise Markov chains [24], provides a foundation for this line of research.
  • 6.
    Study the continuity of the hitting probability map. The discontinuity of the map $T\mapsto p^{T}$ prevented us from adapting the argument for lower hitting probabilities to the upper case. Studying the continuity of this map could allow us to relax the strong conditions that guarantee the uniqueness of $\overline{p}^{\mathcal{T}}$ in Proposition 3.15.
  • 7.
    Relation between lower and upper hitting times and probabilities. Finally, we aim to investigate the relation between hitting times and probabilities in the imprecise framework, and seek to establish whether knowledge of one quantity allows us to infer bounds on the other.
@@ Acknowledgements
This work has been partly supported by the PersOn project (P21-03), which has received funding from Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO). The authors thank the two anonymous reviewers for their careful reading and valuable suggestions that helped improve the exposition. This paper is dedicated to the memory of Prof. Luigi Ferretti, whose kindness and dedication inspired generations of students.
@@ Appendix A Proofs of Auxiliary Lemmas
We now check that $y\in A^{c}\setminus\mathcal{C}_{T}$. First, $y\notin A$: if $y\in A$ then $T(x,y)>0$ would imply $T\mathbb{1}_{A}](x)\geq T(x,y)>0$, contradicting the minimality assumption that $T^{m}\mathbb{1}_{A}](x)=0$ for every $m<n$ (recall $n\geq 2$). Thus $y\in A^{c}$.
Second, $y\notin\mathcal{C}_{T}$: by definition any $z\in\mathcal{C}_{T}$ satisfies $T^{m}\mathbb{1}_{A}](z)=0$ for all $m\in\mathbb{N}$, but we have $T^{n-1}\mathbb{1}_{A}](y)>0$, so $y\notin\mathcal{C}_{T}$. Therefore $y\in A^{c}\setminus\mathcal{C}_{T}$, $T(x,y)>0$ and $T^{n-1}\mathbb{1}_{A}](y)>0$, which proves the lemma. ∎
@@
But $M^{k}\mathbf{0}=\mathbf{0}$ for every $k$. Hence $\mathbf{0}\leq M^{k}\mathbf{1}$. For the upper bound, observe that for every $x\in\mathcal{S}$
@@
Consider any path of length $n_{x}$, $x=x_{0},\dots,x_{n_{x}}$, with $x_{n_{x}}\in A$. Because $m<n_{x}$ implies $T^{m}\mathbb{1}_{A}](x)=0$, such a path cannot visit $A$ before time $n_{x}$. Moreover, by the definition of $\mathcal{C}_{T}$, a path that ever enters $\mathcal{C}_{T}$ cannot reach $A$ at any later time; hence every path contributing to $T^{n_{x}}\mathbb{1}_{A}](x)$ stays in $A^{c}\setminus\mathcal{C}_{T}$} up to time $n_{x}-1$ and then jumps from $A^{c}\setminus\mathcal{C}_{T}$} to $A$ at time $n_{x}$.
@@
Now we use Lemma 2.8: since $M{\mathbf{1}}\leq{\mathbf{1}}$, the map $k\mapsto M^{k}{\mathbf{1}}$ is componentwise nonincreasing: $M^{k+1}{\mathbf{1}}=M^{k}(M{\mathbf{1}})\leq M^{k}{\mathbf{1}}$ for all $k\in\mathbb{N}_{0}$. Let $n\coloneqq\max_{x\in{A^{c}\setminus\mathcal{C}_{T}}}n_{x}$, which is finite because $A^{c}\setminus\mathcal{C}_{T}$} is finite. Then for any $k\geq n$ and any $x\in{A^{c}\setminus\mathcal{C}_{T}}$
$M^{k}{\mathbf{1}}](x)\leq[M^{n_{x}}{\mathbf{1}}](x)<1,$
which is what we wanted to prove. ∎
@@
Let $x\in\mathcal{S}$. Then
$[(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}](x)$ $=\sum_{y\in A^{c}\setminus\mathcal{S}}(T|_{A^{c}\setminus\mathcal{S}})^{k}(x,y)$
$=\ \sum_{\mathclap{\begin{subarray}{c}y\in A^{c}\setminus\mathcal{S}\\ w_{1},\dots,w_{k-1}\in\mathcal{X}\end{subarray}}}\quad T|_{A^{c}\setminus\mathcal{S}}(x,w_{1})T|_{A^{c}\setminus\mathcal{S}}(w_{1},w_{2})\cdots T|_{A^{c}\setminus\mathcal{S}}(w_{k-1},y)$
$=\sum_{\begin{subarray}{c}y\in A^{c}\setminus\mathcal{S}\\ w_{1},\dots,w_{k-1}\in A^{c}\setminus\mathcal{S}\end{subarray}}T(x,w_{1})T(w_{1},w_{2})\cdots T(w_{k-1},y).$
Using $A^{c}\setminus\mathcal{S}\subseteq A^{c}\setminus\mathcal{C}_{T}$ we obtain
$[(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}](x)$ $=\sum_{\begin{subarray}{c}y\in A^{c}\setminus\mathcal{S}\\ w_{1},\dots,w_{k-1}\in A^{c}\setminus\mathcal{S}\end{subarray}}T(x,w_{1})T(w_{1},w_{2})\cdots T(w_{k-1},y)$
$\leq\sum_{\begin{subarray}{c}y\in A^{c}\setminus\mathcal{C}_{T}\\ w_{1},\dots,w_{k-1}\in A^{c}\setminus\mathcal{C}_{T}\end{subarray}}T(x,w_{1})T(w_{1},w_{2})\cdots T(w_{k-1},y)$
$\qquad=[(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}](x)<1,$
which is what we wanted to prove. ∎
@@
Let $\mathcal{S}\subseteq\mathcal{X}$ satisfy $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$ and set